# The Existence of MacWilliams-Type Identities for the Lee, Homogeneous and Subfield Metric

### v2

@@ 3.2. The Homogeneous Metric
Similarly, for any $(I,J)\in\left\{{(S,S),(S,R),(R,S)}\right\}$ we have that the product $a_{k}x_{k}=0$ where $a_{k}\in I$ and $x_{k}\in J$. Therefore, we obtain again
(89) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|.$
@@
Let us now consider $I=U$ the set of units and $a_{k}\in I$. For any $J\in\left\{{U,S,R}\right\}$ we obtain $a_{k}J=J$. Hence, applying Corollary 14, we get
@@
(90) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=\sum_{y\in J}\xi^{\operatorname{Tr}(y)}=\begin{cases}\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle\setminus\left\{{0}\right\}}\xi^{\operatorname{Tr}(y)}=-1&\text{if }J=S,\\ \sum_{y\in\mathcal{R}}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=U,\\ \sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=R.\end{cases}$
@@
Instead, if $J=U$, for $I\in\left\{{Z,U}\right\}$ we already discussed the value of (86) above. In this case, it remains to consider $I=S$ and $I=R$. In the first case, for any $a_{k}\in I$, we have $a_{k}J=I$. However, as the sum runs over $J$, we observe every element in $a_{k}J=I$ exactly $\left|\,{U}\,\right|/\left|\,{S}\,\right|=q^{s-1}$ many times and hence,
(91) $\sum_{x_{k}\in U}\xi^{\operatorname{Tr}(a_{k}x_{k})}=q^{s-1}\sum_{y\in S}\xi^{\operatorname{Tr}(y)}=q^{s-1}\cdot(-1)=-q^{s-1}.$
@@
In the latter case (where $I=R$), consider any $a_{k}\in I$. Then, $a_{k}J=\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle$. As $\sum_{y\in\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0$, we also observe
(92) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(a_{k}x_{k})}=0.$
@@
Lastly, let us consider the case where $I=J=R$. If $a_{k}\in\langle\gamma^{i}\rangle\setminus\langle\gamma^{i+1}\rangle$, then $a_{k}R=(\langle\gamma^{i+1}\rangle\setminus\{0\})^{w}\times(\langle\gamma^{i+1}\rangle)^{v}$, depending on whether the product $a_{k}x_{k}$ can be zero. Note that there are $|\langle\gamma^{s-i}\rangle\setminus\langle\gamma^{s-1}\rangle|=q^{i}-q$ many $x_{k}\in R$ which make the product zero. Thus, the ideal $\langle\gamma^{i+1}\rangle$ of size $q^{s-i-1}$ is seen $v=q^{i}-q$ many times and we are left with $|R|-(q^{i}-q)q^{s-i-1}=q^{s-i}-q$ many $x_{k}\in R$, which see $\langle\gamma^{i+1}\rangle\setminus\{0\}$ exactly $w=(q^{s-i}-q)/(q^{s-i-1}-1)=q$ times. Thus, we get that
(93) $\sum_{x_{k}\in R}\xi^{\operatorname{Tr}{(x_{k}a_{k})}}=(q^{i}-q)\sum_{y\in\langle\gamma^{i+1}\rangle}\xi^{\operatorname{Tr}{(y)}}+q\sum_{y\in\langle\gamma^{i+1}\rangle\setminus\{0\}}\xi^{\operatorname{Tr}{(y)}}=(q^{i}-q)\cdot 0+q\cdot(-1)=-q.$
Table 1 gives an overview of the values of the sum in (86) for all $I,J\in\left\{{Z,U,S,R}\right\}$. It is worth noting that these values are independent of the choice of $a_{k}\in I$. This independence yields to a well-defined expression for the Krawtchouk coefficients, wherefore they exist. By repeating the procedure for each position $k$ and summing over all possible compositions $t\in\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)$, we obtain the desired result.
Table 1. Values of the expression $\sum_{x_{\ell}\in J}\xi^{\operatorname{Tr}(a_{k}\cdot x_{\ell})}$ for any $I,J\in\left\{{Z,U,S,R}\right\}$ as given in (86).
$I$ $J$ $Z$ $U$ $S$ $R$
$Z$ 1 $q^{s-1}(q-1)$ $q-1$ $q^{s-1}-q$
$U$ 1 0 $-1$ 0
$S$ 1 $-q^{s-1}$ $q-1$ $q^{s-1}-q$
$R$ 1 0 $q-1$ $-q$
Let us consider an example to get an idea of the formula stated above.
@@
Consider again the code $\mathcal{C}\subset(\mathbb{Z}/{9}\mathbb{Z})^{3}$ generated by $G=\begin{pmatrix}3&2&8\end{pmatrix}$. Notice, that there are three distinct homogeneous weight-unit decompositions occurring in $\mathcal{C}$, namely,
(94) $\pi^{(0)}=(3,0,0,0),\quad\pi^{(1)}=(1,0,2,0),\quad\pi^{(2)}=(0,2,1,0),$
(95) $\mathcal{D}_{\pi^{(0)}}^{\mathsf{Hom}}(\mathcal{C})=1,\quad\mathcal{D}_{\pi^{(1)}}^{\mathsf{Hom}}(\mathcal{C})=2,\quad\mathcal{D}_{\pi^{(2)}}^{\mathsf{Hom}}(\mathcal{C})=6.$
Note that the dual $\mathcal{C}^{\perp}$ of $\mathcal{C}$ is generated by
(96) $G^{\perp}=\begin{pmatrix}1&0&3\\ 0&1&2\end{pmatrix}.$
Let $\rho=(2,0,1,0)\in\mathbb{D}_{9,3}^{\mathsf{Hom}}$ be a homogeneous weight-unit decomposition in $(\mathbb{Z}/{9}\mathbb{Z})^{3}$. By calculating the codewords of $\mathcal{C}^{\perp}$, we observe that $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})=2$. For every $\pi\in\left\{{\pi^{(0)},\pi^{(1)},\pi^{(2)}}\right\}$ let us now compute the Krawtchouk coefficient $K^{\mathsf{Hom}}_{\pi}(\rho)$ as stated in Theorem 30.
(97) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)=\Big{\{}(\underbrace{2,0,1,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{0,0,0,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}})\Big{\}}.$
Hence, we obtain
(98) $K^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)=\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}=\binom{\pi_{Z}}{t_{Z}}(p-1)^{t_{ZS}}=\binom{3}{2,0,1,0}2^{1}=6.$
Now let $\pi=\pi^{(1)}$. Note that the set of compositions $\mathrm{Comp}^{\mathsf{Hom}}_{\pi}(\rho)$ of $\rho$ with respect to $\pi$ consists of only two elements, i.e.,
(99) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)=\Big{\{}t=(\underbrace{1,0,0,0}_{t_{Z}},\underbrace{0,0,0,0}_{t_{U}},\underbrace{1,0,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),s=(\underbrace{0,0,1,0}_{s_{Z}},\underbrace{0,0,0,0}_{s_{U}},\underbrace{2,0,0,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big{\}}.$
With this we obtain
(100) $K^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)$ $=\left(\binom{\pi_{Z}}{t_{Z}}\binom{\pi_{S}}{t_{S}}(p-1)^{t_{SS}}\right)+\left(\binom{\pi_{Z}}{s_{Z}}\binom{\pi_{S}}{s_{S}}(p-1)^{s_{SS}}\right)$
(101) $=\binom{1}{1,0,0,0}\binom{2}{1,0,1,0}2^{1}+\binom{1}{0,0,1,0}\binom{2}{2,0,0,0}2^{1}$
(102) $=6.$
Lastly, for $\pi^{(2)}$ we have
(103) $\mathrm{Comp}^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)=\Big{\{}t=(\underbrace{0,0,0,0}_{t_{Z}},\underbrace{2,0,0,0}_{t_{U}},\underbrace{0,0,1,0}_{t_{S}},\underbrace{0,0,0,0}_{t_{R}}),\ s=(\underbrace{0,0,0,0}_{s_{Z}},\underbrace{1,0,1,0}_{s_{U}},\underbrace{1,0,0,0}_{s_{S}},\underbrace{0,0,0,0}_{s_{R}})\Big{\}},$
which yields to
(104) $K^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)$ $=\left(\binom{\pi_{U}}{t_{U}}\binom{\pi_{S}}{t_{S}}(p-1)^{t_{SS}}\right)+\left(\binom{\pi_{U}}{s_{U}}\binom{\pi_{S}}{s_{S}}(-1)^{s_{US}}\right)$
(105) $=\binom{2}{2,0,0,0}\binom{1}{0,0,1,0}2^{1}+\binom{2}{1,0,1,0}\binom{1}{1,0,0,0}(-1)^{1}$
(106) $=0.$
Finally, we apply Equation (82) and we indeed obtain
(107) $\mathcal{D}_{\rho}^{\mathsf{Hom}}(\mathcal{C}^{\perp})$ $=\frac{1}{9}\left(K^{\mathsf{Hom}}_{\pi^{(0)}}(\rho)\mathcal{D}_{\pi^{(0)}}^{\mathsf{Hom}}+K^{\mathsf{Hom}}_{\pi^{(1)}}(\rho)\mathcal{D}_{\pi^{(1)}}^{\mathsf{Hom}}+K^{\mathsf{Hom}}_{\pi^{(2)}}(\rho)\mathcal{D}_{\pi^{(2)}}^{\mathsf{Hom}}\right)$
(108) $=\frac{1}{9}(6\cdot 1+6\cdot 2+0\cdot 6)=2.$
Finally, note that the homogeneous unit-weight partition (see Definition 29) defines the coarsest partition that allows us to obtain MacWilliams-type identities for the homogeneous metric. In fact, we are looking for a partition of the code that gives well-defined Krawtchouk coefficients and is finer than the homogeneous-weight partition, as we want to recover the weight enumerator.
3.3. The Subfield Metric
Let us focus on the subfield metric defined over the finite field $\mathbb{F}_{p^{r}}$. Since $\mathbb{F}_{p^{r}}$ is a finite chain ring with $s=1$ and the subfield weight is additive, using the Lee partition allows us to obtain MacWilliams-type identities for the subfield weight. However, this is not the coarsest partition we can choose. Again, a natural choice is to consider the subfield decomposition of the codewords defined as follows: given a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ the subfield decomposition of a codeword $c\in\mathcal{C}$ is a triple $(s_{0}(c),s_{1}(c),s_{\lambda}(c))$, where for each $i\in\left\{{0,1,\lambda}\right\}$ we define
(109) $s_{i}(c):=\left|\,{\left\{{j=1,\ldots,n\;|\;\operatorname{wt}_{\lambda}(c_{j})=i}\right\}}\,\right|.$
We then can partition the code into blocks containing the codewords with the same subfield weight decomposition. The sizes of the blocks determine the subfield decomposition enumerator of a code $\mathcal{C}$ and we denote it by $\lambda\mathsf{dec}_{\mathcal{C}}$.
Example 32.
@@ 3.3. The Subfield Metric
Consider the finite field $\mathbb{F}_{8}=\mathbb{F}_{2}[\alpha]$ where $\alpha^{3}=\alpha+1$ and the two codes $\mathcal{C}_{1}=\left\langle{\begin{pmatrix}1,\alpha,\alpha^{2}\end{pmatrix}}\right\rangle$ and $\mathcal{C}_{2}=\left\langle{\begin{pmatrix}\alpha^{2}+2,2\alpha+1,\alpha^{2}+1\end{pmatrix}}\right\rangle$. We denote by $\mathcal{C}^{\perp}_{1}$ and $\mathcal{C}^{\perp}_{2}$ the dual of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Table 2 then shows that $\lambda\mathsf{dec}_{\mathcal{C}_{1}}=\lambda\mathsf{dec}_{\mathcal{C}_{2}}$, but $\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{2}}$.
@@
Note that Example 32 also shows that partitioning a code $\mathcal{C}$ into codewords of the same subfield weight does not yield to MacWilliams identities. Indeed, for $\lambda>1$, Example 32 shows that the subfield weight enumerators of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ coincide (i.e., $\lambda\mathsf{we}_{\mathcal{C}_{1}}=\lambda\mathsf{we}_{\mathcal{C}_{2}}$), whereas the subfield weight enumerators of their duals are different (i.e, $\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{2}}$).
Remark 33.
For the MacWilliams-type identities for the subfield weight we consider a character of the form $\chi_{a}(x)=\xi^{\operatorname{Tr}(\left\langle{a,x}\right\rangle)}$ of $\mathbb{F}_{p^{r}}^{n}$. For quantum error-correcting codes, the use of the alternating or Hermitian form instead of the standard inner product are usually a more natural choice (see [8, 22] for more details).
However, the Hermitian form for any two vectors $x,y\in\mathbb{F}_{p^{r}}^{n}$ is defined for $r=2$, for which the MacWilliams identities have been shown to exist [22]. The alternating inner product of two vectors $x,y\in\mathbb{F}_{p^{r}}^{n}$ is defined only for an even $r$. We fix a $\mathbb{F}_{p}$-basis $\Gamma=\{1,\gamma,\ldots,\gamma^{r-1}\}$ of $\mathbb{F}_{p^{r}}$ and expand the vectors over the base field. That is,for any coordinate $1\leq i\leq n$
(110) $x_{i}=x_{i}^{(0)}+x_{i}^{(1)}\gamma+\ldots+x_{i}^{(r-1)}\gamma^{r-1}\in\mathbb{F}_{p^{r}}$
and we write $\Gamma(x_{i})=(x_{i}^{(0)},\ldots,x_{i}^{(s-1)})\in\mathbb{F}_{p^{r}}.$ Then, we denote by $x\star y=\sum_{i=1}^{n}x_{i}\star y_{i}$ the alternating inner product of $x,y\in\mathbb{F}_{p^{r}}$ where
(111) $x_{i}\star y_{i}=\sum_{j=0}^{r/2-1}x_{i}^{(j)}y_{i}^{(j+r/2)}-x_{i}^{(j+r/2)}y_{i}^{(j)}.$
Then a character of $\mathbb{F}_{p^{r}}$ has the form $\chi_{a}(x)=\xi^{\operatorname{Tr}(a\star x)}$. Notice that finding a partition $\mathcal{P}=P_{1}\;|\;\ldots\;|\;P_{t}$ of $\mathbb{F}_{p^{r}}^{n}$ that allows for an expression of $\sum_{x\in P_{j}}\xi^{\operatorname{Tr}(a\star x)}$ that is independent on the choice of $a\in P_{i}$ is way more involved than for the case where we consider the standard inner product. In fact, to get a well-defined Krawtchouk coefficient one would need to go through all possible $\Gamma(x),$ leaving us with $p^{rn}$ choices.
Another idea, similar to the Lee partition, is to consider an orbit partition for which, by [19], MacWilliams identities hold. Based on the definition of the subfield weight over $\mathbb{F}_{p^{r}}$, we consider the multiplicative group action of $\mathbb{F}_{p}^{\times}$ on $\mathbb{F}_{p^{r}}$. We partition the finite field $\mathbb{F}_{p^{r}}$ into distinct orbits of multiplication with $\mathbb{F}_{p}^{\times}$, i.e., $I_{\alpha}:=\alpha\mathbb{F}_{p}^{\times}$, where $\alpha\in\mathbb{F}_{p^{r}}$. To keep elements of different subfield weight separate, it is natural to consider the orbits $I_{0}:=\left\{{0}\right\}$ and $I_{1}:=\mathbb{F}_{p}^{\times}$. In the following we denote $\alpha_{0}:=0$ and $\alpha_{1}:=1$ and write $I_{0}=\alpha_{0}\mathbb{F}_{p}^{\times}$ and $I_{1}=\alpha_{1}\mathbb{F}_{p}^{\times}$. Then, for $i=2,\ldots,\frac{p^{r}-1}{p-1}$, we recursively choose an element $\alpha_{i}\in\mathbb{F}_{p^{r}}\setminus\left(\bigcup_{j=0}^{i-1}I_{j}\right)$ and define $I_{i}:=\alpha_{i}\mathbb{F}_{p}^{\times}.$ Defining $\ell:=\frac{p^{r}-1}{p-1}$, we get that
(112) $\mathcal{I}=I_{0}\;|\;\ldots\;|\;I_{\ell},$
forms a partition of $\mathbb{F}_{p^{r}}$ which we call the subfield-orbit partition of $\mathbb{F}_{p^{r}}$. The subfield-orbit partition $\mathcal{I}^{\lambda}$ allows to define the subfield-orbit decomposition of a tuple $x\in\mathbb{F}_{p^{r}}^{n}$. That is, given a vector $x\in\mathbb{F}_{p^{r}}^{n}$ we define its subfield-orbit decomposition to be the $(\ell+1)$-tuple $\sigma(x)=(\sigma_{0}(x),\sigma_{1}(x),\ldots,\sigma_{\ell}(x))$ where, for every $i\in\left\{{0,\ldots,\ell}\right\}$,
(113) $\sigma_{i}(x):=\left|\,{\left\{{k=1,\ldots,n\;|\;x_{k}\in I_{i}}\right\}}\,\right|.$
Given a linear code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$, we can partition the codewords of $\mathcal{C}$ via their subfield-orbit partition. This is a partition of $\mathcal{C}$ given by the orbits of $\mathbb{F}_{p}^{\times}$ which form a subgroup of the automorphism group of $\mathbb{F}_{p^{r}}$ and, hence, we obtain MacWilliams identities for the subfield-orbit enumerator of the code $\mathcal{C}$. Additionally, note that every $c\in P^{\lambda}_{\pi}$ has the same subfield weight given by $\operatorname{wt}_{\lambda}(c)=\sigma_{1}+\lambda\sum_{i=2}^{\ell}\sigma_{i}$. Hence, the subfield weight enumerator $\lambda\mathsf{we}_{\mathcal{C}}$ can immediately be derived from the subfield-orbit enumerator.
@@
The subfield-orbit partition of $\mathbb{F}_{p^{r}}^{n}$, however, is not the coarsest partition that recovers the subfield weight enumerator and allows for MacWilliams-type identities. In fact, we can do better by defining a reflexive partition consisting of only four blocks that is not based on the orbits of a subgroup. More explicitly, with an eye on Example 11, we will distinguish the elements in $\mathbb{F}_{p^{r}}^{n}$ not only by their subfield weight, but also by their trace. As we still want to be able to recover the subfield weight of a vector, we subpartition the elements of subfield weight $\lambda$ (i.e., the elements in $\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}$) into two sets: the first one contains all such elements with trace equal to 00 and the second one contains all remaining elements. That is, we use $\mathcal{I}^{\lambda}=I^{\lambda}_{0}\;|\;I^{\lambda}_{1}\;|\;I^{\lambda}_{2}\;|\;I^{\lambda}_{3}$ to partition $\mathbb{F}_{p^{r}}$ where

### v3

@@ 3.2. Symmetrized Homogeneous-Unit Partition
where the cardinalities $\left|\,{J}\,\right|$ with $J\in\left\{{Z,U,S,R}\right\}$ are given in (49). Similarly, for any $(I,J)\in\left\{{(S,S),(S,R),(R,S)}\right\}$ we have that the product $ax=0$ where $a\in I$ and $x\in J$. Therefore, we obtain again
(55) $\sum_{x_{k}\in J}\xi^{\operatorname{Tr}(ax_{k})}=\sum_{y\in J}\xi^{0}=\left|\,{J}\,\right|.$
@@
Let us now consider $I=U$ the set of units and $a\in I$. For any $J\in\left\{{U,S,R}\right\}$ we obtain $aJ=J$. Hence, applying Corollary 13, we get
@@
(56) $\sum_{x\in J}\xi^{\operatorname{Tr}(ax)}=\begin{cases}\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle\setminus\left\{{0}\right\}}\xi^{\operatorname{Tr}(y)}=-1&\text{if }J=S,\\ \sum_{y\in\mathcal{R}}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=U,\\ \sum_{y\in\left\langle{\gamma}\right\rangle}\xi^{\operatorname{Tr}(y)}-\sum_{y\in\left\langle{\gamma^{s-1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0&\text{if }J=R.\end{cases}$
@@
Instead, if $J=U$, it remains to consider the cases $I=S$ and $I=R$. In the first case, for any $a\in I$, we have $aJ=I$. However, as the sum runs over $J$, we observe every element in $aJ=I$ exactly $\left|\,{U}\,\right|/\left|\,{S}\,\right|=q^{s-1}$ many times and hence,
(57) $\sum_{x\in U}\xi^{\operatorname{Tr}(ax)}=q^{s-1}\sum_{y\in S}\xi^{\operatorname{Tr}(y)}=q^{s-1}\cdot(-1)=-q^{s-1}.$
@@
In the latter case $I=R$, consider any $a\in I$. Then, $aJ=\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle$. Since $\sum_{y\in\left\langle{\gamma^{i}}\right\rangle\setminus\left\langle{\gamma^{i+1}}\right\rangle}\xi^{\operatorname{Tr}(y)}=0$, we also observe
(58) $\sum_{x\in J}\xi^{\operatorname{Tr}(ax)}=0.$
@@
Lastly, let us consider the case where $I=J=R$. If $a\in\langle\gamma^{i}\rangle\setminus\langle\gamma^{i+1}\rangle$, then $aR=(\langle\gamma^{i+1}\rangle\setminus\{0\})^{w}\times(\langle\gamma^{i+1}\rangle)^{v}$, for some positive integers $v$ and $w$, depending on whether the product $ax$ can be zero. Note that there are $|\langle\gamma^{s-i}\rangle\setminus\langle\gamma^{s-1}\rangle|=q^{i}-q$ many $x\in R$ which make the product zero. Thus, the ideal $\langle\gamma^{i+1}\rangle$ of size $q^{s-i-1}$ is seen $v=q^{i}-q$ many times and we are left with $|R|-(q^{i}-q)q^{s-i-1}=q^{s-i}-q$ many $x\in R$, which see $\langle\gamma^{i+1}\rangle\setminus\{0\}$ exactly $w=(q^{s-i}-q)/(q^{s-i-1}-1)=q$ times. Thus, we get that
(59) $\sum_{x\in R}\xi^{\operatorname{Tr}{(xa)}}$ $=(q^{i}-q)\sum_{y\in\langle\gamma^{i+1}\rangle}\xi^{\operatorname{Tr}{(y)}}+q\sum_{y\in\langle\gamma^{i+1}\rangle\setminus\{0\}}\xi^{\operatorname{Tr}{(y)}}$
(60) $=(q^{i}-q)\cdot 0+q\cdot(-1)$
(61) $=-q.$
Table 1 gives an overview of the values of the sum $\sum_{x\in J}\chi_{a}(x)$ for all $I,J\in\left\{{Z,U,S,R}\right\}$ and $a\in I$. The rows of the table are indexed by the blocks of $\widehat{\mathcal{P}^{\mathsf{hom}}}$ and the columns by the blocks of $\mathcal{P}^{\mathsf{hom}}$, showing that $\mathcal{P}^{\mathsf{hom}}$ and $\widehat{\mathcal{P}^{\mathsf{hom}}}$ contain the same number of blocks, which completes the proof.
Table 1. Values of the expression $\sum_{x\in J}\xi^{\operatorname{Tr}(a\cdot x)}$ for any $I,J\in\left\{{Z,U,S,R}\right\}$.
II $J$ $Z$ $U$ $S$ $R$
$Z$ 1 $q^{s-1}(q-1)$ q1q-1 $q^{s-1}-q$
$U$ 1 0 1-1 0
$S$ 1 $q^{s-1}$ q1q-1 $q^{s-1}-q$
$R$ 1 0 q1q-1 q-q
By [18, Theorem 3.5] the MacWilliams-type identities exist for the symmetrized homogeneous weight-unit enumerator and can be summarized as follows.
Corollary 28 (MacWilliams-type Identities for the Symmetrized Homogeneous Weight-unit Enumerator).
Let $\mathcal{R}$ be a finite chain ring with residue field size $q=p^{r}$ and nilpotency index $s$. Consider a linear code $\mathcal{C}\subseteq\mathcal{R}^{n}$ and let $\mathcal{C}^{\perp}$ denote its dual. For a given homogeneous weight-unit decomposition $\rho\in\mathbb{D}^{\mathsf{hom}}_{n}$, the number of codewords in the dual $\mathcal{C}^{\perp}$ of homogeneous weight-unit decomposition $\rho$ is given by
(62) $\mathcal{D}_{\rho}^{\mathsf{hom}}(\mathcal{C}^{\perp})=\frac{1}{\left|\,{\mathcal{C}}\,\right|}\sum_{\pi\in\mathbb{D}^{\mathsf{hom}}_{n}}K^{\mathsf{hom}}_{\pi}(\rho)\mathcal{D}_{\pi}^{\mathsf{hom}}(\mathcal{C}).$
Here, $K^{\mathsf{hom}}_{\pi}(\rho)$ denotes the value of the Krawtchouk coefficient of $\rho$ with respect to $\pi$, and for a given $p^{s}$-th root of unity $\xi$ is given by
(63) $K^{\mathsf{hom}}_{\pi}(\rho)=\sum_{t\in\mathrm{Comp}^{\mathsf{hom}}_{\pi}(\rho)}\Bigg($ $\prod_{i\in\left\{{Z,U,S,R}\right\}}\binom{\pi_{i}}{t_{iZ},\ldots,t_{iR}}\Bigg)\kappa^{\mathsf{hom}},$
where, given that $t_{UU}=t_{UR}=t_{RU}=0$, we have
(64) $\kappa^{\mathsf{hom}}=$ $(-1)^{t_{US}}(-q^{s-1})^{t_{SU}}\left(q^{s-1}(q-1)\right)^{t_{ZU}}\left(q^{s-1}-q\right)^{t_{ZR}}(q-1)^{t_{ZS}+t_{SS}+t_{SR}+t_{RS}}(-q)^{t_{RR}}.$
The proof follows a similar arguments as the proof of Corollary 24.
In A.2, we give an example to get an idea of the formula stated above. Finally, note that the homogeneous unit-weight partition defines the coarsest partition that allows us to obtain MacWilliams-type identities for the homogeneous metric.
Lemma 29.
Given a finite chain ring $\mathcal{R}$ and a code $\mathcal{C}\subseteq\mathcal{R}^{n}$. The symmetrized homogeneous weight-unit partition $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{hom}}$ is the coarsest symmetrized partition which is Fourier-reflexive and allows to recover the homogeneous weight enumerator $\mathcal{C}$.
Proof.
Let us assume there exists partition $\mathcal{Q}_{\text{sym},n}$ of $\mathcal{R}^{n}$ which is coarser than $\mathcal{P}_{\mathrm{sym},n}^{\mathsf{hom}}$. Hence, the underlying partition $\mathcal{Q}$ of $\mathcal{R}$ consists of $3$ or fewer blocks. If $\mathcal{Q}$ is not the homogeneous weight partition, by similar arguments as in the proof of Lemma 25, the homogeneous weight enumerator cannot be recovered. On the other hand, if $\mathcal{Q}$ is the homogeneous weight partition, then $\mathcal{Q}_{\text{sym},n}$ is not Fourier-reflexive in general (see Theorem [26]). ∎
3.3. Symmetrized Subfield-Trace Partition
Let us focus on the subfield metric defined over the finite field $\mathbb{F}_{p^{r}}$. Since $\mathbb{F}_{p^{r}}$ is a finite chain ring with $s=1$ and the subfield weight is additive, the symmetrized Lee partition allows us to obtain MacWilliams-type identities for the subfield weight. However, this is not the coarsest partition we can choose. Again, a natural choice is to consider the subfield decomposition of the codewords defined as follows: given a code $\mathcal{C}\subseteq\mathbb{F}_{p^{r}}^{n}$ the subfield decomposition of a codeword $c\in\mathcal{C}$ is a triple $(s_{0}(c),s_{1}(c),s_{\lambda}(c))$, where for each $i\in\left\{{0,1,\lambda}\right\}$ we define
(65) $s_{i}(c):=\left|\,{\left\{{j\in\{1,\ldots,n\}\;|\;\operatorname{wt}_{\lambda}(c_{j})=i}\right\}}\,\right|.$
We can then partition the code into blocks containing the codewords with the same subfield weight decomposition. The sizes of the blocks determine the symmetrized subfield decomposition enumerator of a code $\mathcal{C}$ and we denote it by $\lambda\mathsf{dec}_{\mathcal{C}}$.
@@ 3.3. Symmetrized Subfield-Trace Partition
Consider the finite field $\mathbb{F}_{8}=\mathbb{F}_{2}(\alpha)$ where $\alpha^{3}=\alpha+1$ and the two codes $\mathcal{C}_{1}=\left\langle{\begin{pmatrix}1,\alpha,\alpha^{2}\end{pmatrix}}\right\rangle$ and $\mathcal{C}_{2}=\left\langle{\begin{pmatrix}\alpha^{2},1,\alpha^{2}+1\end{pmatrix}}\right\rangle$. We denote by $\mathcal{C}^{\perp}_{1}$ and $\mathcal{C}^{\perp}_{2}$ the dual of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$, respectively. Table 2 then shows that $\lambda\mathsf{dec}_{\mathcal{C}_{1}}=\lambda\mathsf{dec}_{\mathcal{C}_{2}}$, but $\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{dec}_{\mathcal{C}^{\perp}_{2}}$.
@@
Example 30 also shows that partitioning a code $\mathcal{C}$ into codewords of the same subfield weight does not yield to MacWilliams identities. Indeed, for $\lambda>1$, Example 30 shows that the subfield weight enumerators of $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ coincide (i.e., $\lambda\mathsf{we}_{\mathcal{C}_{1}}=\lambda\mathsf{we}_{\mathcal{C}_{2}}$), whereas the subfield weight enumerators of their duals are different (i.e, $\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{1}}\neq\lambda\mathsf{we}_{\mathcal{C}^{\perp}_{2}}$).
@@
The symmetrized subfield decomposition defined above is the coarsest partition that allows to recover the subfield weight of the elements of a code. As seen in Example 30, it is not Fourier-reflexive. However, we can define a Fourier-reflexive partition consisting of only four blocks. More explicitly, with an eye on Example 9, we will distinguish the elements in $\mathbb{F}_{p^{r}}^{n}$ not only by their subfield weight, but also by their trace. As we still want to be able to recover the subfield weight of a vector, we sub-partition the elements of subfield weight $\lambda$ (i.e., the elements in $\mathbb{F}_{p^{r}}\setminus\mathbb{F}_{p}$) into two sets: the first one contains all such elements with trace equal to 0 and the second one contains all remaining elements. That is, we partition $\mathbb{F}_{p^{r}}$ as $\mathcal{P}^{\lambda}=I^{\lambda}_{0}\;|\;I^{\lambda}_{1}\;|\;I^{\lambda}_{2}\;|\;I^{\lambda}_{3}$, where